The length of a rectangle is fixed at 10cm. What widths will make the perimeter greater than 88cm?
The width must be greater than _______ cm.
: The length of a rectangle is fixed at 10cm. What widths will make the perimeter greater than 88cm?
The width must be greater than _______ cm.

Answer by oscargut(667)

Perimeter is p=2(10+w)=20+2w
and i want p>88 so 20+2w>88
2w>68
w>34
The width must be greater than __34_____ cm.